Three students, Thando, Busi and Peter, painted a room together. They spent 5 hours, 3 hours and 2 hour, respectively, on the task. They earned R500 in total, for the job. How much should each of them receive?
1
Expert's answer
2019-08-15T10:01:51-0400

The three students spent a total of "5+2+3=10 hrs"

Thando should receive "\frac{5}{10}\times{500}=R 250"

Busi should receive "\frac{3}{10}\times{500}=R 150"

Peter should receive "\frac{2}{10}\times{500}=R 100"